# Geometry of polynomial neural networks

### v1

@@ 1. Introduction
In particular, researchers study various activation functions to understand their role in introducing non-linearity, affecting gradient propagation, and influencing computational efficiency. Polynomial activation functions have gained interest for their ability to introduce higher-order interactions between inputs, allowing networks to model complex, non-linear phenomena more effectively [OPP03]. Although feedforward neural networks with polynomial activation functions are well-known to be non-universal approximators [HSW89], there are many practical tasks where architectures with polynomial activations have outperformed other ones, especially in environments where data relationships are polynomial in nature. In particular, polynomial neural networks have led to state-of-the-art results in engineering tasks like face detection from cluttered images [HSHK03], image generation [${}^{+}$20], 3D shape recognition [${}^{+}$21], as well as financial applications like forecasting trading signals [GHL11], uncertain natural frequency quantification [${}^{+}$16], and estimating stock closing indices [NM18]. However, the choice of degree and the potential for overfitting are critical considerations in applying polynomial neural networks to individual practical tasks.
@@
In this paper, we perform an algebro-geometric study of neuromanifolds and their Zariski closures, neurovarieties, following the previous work by [KTB19]. In §2, we give the background on polynomial neural networks, neuromanifolds and neurovarieties. Neuromanifolds provide a natural generalization of the set of symmetric tensors of bounded symmetric rank. In §3, we recall the connection between homogeneous polynomials and symmetric tensor decompositions, and characterize neuromanifolds for some shallow polynomial neural networks. In §4, we describe different approaches for studying neurovarieties. The dimension of the neurovariety provides a measure for the expressivity of the network. In §5, we study these dimensions. For a shallow polynomial neural network with a single output unit, this corresponds to the Alexander–Hirschowitz Theorem in neural network settings. In the deep case, we present conjectures supported by empirical evidence. Finally, in §6, we study the optimization process of a polynomial neural network: from a static perspective, we describe the complexity of the optimization landscape by its learning degree and compute it for a family of architectures; for the dynamic optimization process, we review the backpropagation algorithm and summarize known results on linear neural networks from previous literature in both the machine learning and algebraic statistics communities.
@@
Acknowledgments. This paper was written in connection with the “Apprenticeship Week: Varieties from Statistics” which took place at IMSI as part of the long program “Algebraic Statistics and Our Changing World”. We would like to thank Miles Bakenhus and Maksym Zubkov who participated in the work of this group part of the time. Lemma 3.5 is due to Maksym Zubkov. We would like to thank Elizabeth Gross, Leonie Kayser, Joe Kileel, Mark Kong, Guido Montúfar, Alessandro Oneto, Bernd Sturmfels, and Matthew Trager for useful discussions and various suggestions. We thank Lisa Seccia and Nils Sturma for careful reviews of an earlier version of this paper. Part of this research was performed while the authors were visiting the Institute for Mathematical and Statistical Innovation (IMSI), which is supported by the National Science Foundation (Grant No. DMS-1929348).
@@ 2. Background
Figure 1. A neural network architecture with widths $\mathbf{d}=(2,3,3,2)$, input $\mathbf{x}=(x_{1},x_{2})^{T}$ and output $y_{1}$.
Pictorially, a neural network architecture can be represented as in Figure 1. Each node is called a neuron and each column of neurons forms a layer of the network. The first layer is called input layer, the last layer is called output layer and the remaining layers are referred to as hidden layers. A neural network with exactly one hidden layer is called shallow. The number of neurons in the $i^{\text{th}}$ layer is the $i^{\text{th}}$ width, denoted $d_{i}$. The vector of widths $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ together with a choice of activation functions $\bm{\sigma}=(\sigma_{1},\sigma_{2},\dots,\sigma_{L-1})$ constitute the architecture of the network. Arrows between layers represent the maps $f_{l}$.
@@
The image of $\Psi_{\mathbf{d},r}$ is the neuromanifold $\mathcal{M}_{\mathbf{d},r}$; it is a semialgebraic set inside $(\operatorname{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$. Its Zariski closure is the neurovariety $\mathcal{V}_{\mathbf{d},r}$, an affine variety inside $(\operatorname{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$.
@@
[…]
@@
Most ideals of neurovarieties in Table 1 are described by results in the upcoming sections. The ideals for all but two cases ($\mathbf{d}=(3,2,3)$ and $\mathbf{d}=(3,3,3)$) in the table are determinantal. Only the neurovariety for the architecture $\mathbf{d}=(3,2,1)$ does not have expected dimension. The following example describes the cases that are not explained by any results in the rest of the paper.
@@ 3.3. Neuromanifolds for $r\in\mathbb{N}$
Let $S^{n}_{d,k}=\{T\in\operatorname{Sym}_{d}(\mathbb{R}^{n}):\text{ symmetric rankof }T\text{ is }k\}$; if $S^{n}_{d,k}\subseteq\operatorname{Sym}_{d}(\mathbb{R}^{n})$ has non-empty interior with respect to the Euclidean topology, $k$ is called a typical symmetric rank for order-$d$ tensors of dimension $n$.
Over the complex numbers, there is a unique typical symmetric rank which is called the generic symmetric rank. However, over the real numbers there can be multiple typical symmetric ranks. A neuromanifold $\mathcal{M}_{\mathbf{d},r}$ having non-empty interior with respect to the Euclidean topology implies that the architecture $(\mathbf{d},r)$ is filling. The importance of typical symmetric ranks for PNNs comes from the following fact which is a consequence of the discussions above and the inclusion $S^{n}_{d,k}\subseteq\mathcal{M}_{(n,k,1),d}$.
@@
Little is known about the algebraic boundaries in these cases. Theorem 4.1 in [MMSV16] implies that $\partial_{\text{alg}}(\mathcal{M}_{(3,6,1),4})$ contains an irreducible component of degree 51.
@@ 4. Neurovarieties
[…]
@@
$\begin{pmatrix}[r]c_{1112}^{(2)}&-c_{1111}^{(2)}&\makebox[0.0pt][c]{0}&c_{1112}^{(1)}&-c_{1111}^{(1)}&\makebox[0.0pt][c]{0}\\ 4c_{1122}^{(2)}&-c_{1112}^{(2)}&8c_{1111}^{(2)}&4c_{1122}^{(1)}&-c_{1112}^{(1)}&8c_{1111}^{(1)}\\ 8c_{1222}^{(2)}&\makebox[0.0pt][c]{0}&8c_{1112}^{(2)}&8c_{1222}^{(1)}&\makebox[0.0pt][c]{0}&8c_{1112}^{(1)}\\ 8c_{2222}^{(2)}&c_{1222}^{(2)}&4c_{1122}^{(2)}&8c_{2222}^{(1)}&c_{1222}^{(1)}&4c_{1122}^{(1)}\\ \makebox[0.0pt][c]{0}&c_{2222}^{(2)}&c_{1222}^{(2)}&\makebox[0.0pt][c]{0}&c_{2222}^{(1)}&c_{1222}^{(1)}\\ \end{pmatrix}.$
@@ 5.1. A plethora of conjectures
It is conjectured that the neurovariety attains the expected dimension for large activation degree. Compare this for example with the Alexander–Hirschowitz Theorem which tells us there exist no defective single-output two-layer neurovarieties with activation degree $r\geq 5$.
@@ 6. Optimization
In order to minimize the empirical risk, one typically uses gradient-based optimization methods, such as gradient descent (GD), stochastic gradient descent (SGD), and adaptive moment estimation (Adam). With the proper choice of step size, gradient-based algorithms are guaranteed to converge to the global minimum when the objective function is convex. However, when optimizing over a neural network, the loss landscape is often highly non-convex regardless of the choice of $L$. None of the gradient-based algorithms mentioned above are guaranteed to find the optimal $f^{*}$ for general neural networks.
There have been two lines of work studying the optimization process of neural networks: One can characterize the geometry of the loss landscape, including the dimension, critical points, local and global minima; we refer to these as the static properties. From another perspective, one can also track the trajectory of the optimization algorithms, understanding the solution an algorithm converges to with respect to different initialization; we refer to this as the dynamic process.
@@ 6.1. The Learning Degree of a PNN
Let us consider the case of PNNs again. Often one tries to solve the optimization problem (6.1) using gradient based methods, see §6.2. Therefore, one is interested in computing the critical points on the neurovariety of the loss function $\ell_{p_{\mathbf{w}}}$. In particular, the number of such points provides an upper bound for the number of functions the optimization process can converge to when using gradient descent.
If there is a finite number of critical points of $\ell_{p_{\mathbf{w}}}$ which is constant for a general choice of training data $(\mathbf{x}_{i},\mathbf{y}_{i})_{i}$, we call this number the learning degree of the network $p_{\mathbf{w}}$ with respect to loss $\ell$.
In the following we are studying the case of the Euclidean loss function. We show that in this case the learning degree is independent of the choice of training data $(\mathbf{x}_{i},\mathbf{y}_{i})_{i}$ and is in fact equal to the generic Euclidean distance (ED) degree. We would like to thank Matthew Trager for pointing out the connection to the generic ED degree.
@@
Let $(p_{\mathbf{w}}^{(i)}(\mathbf{x}))_{i}\in(\operatorname{Sym}_{d}(\mathbb{R}^{n}))^{m}$ be a tuple of polynomials in $n$ variables $\mathbf{x}=(x_{1},\dots,x_{n})$ occurring as the output of a fixed PNN with weight vector $\mathbf{w}$. Suppose we want the PNN to learn the tuple of polynomials $(f_{i}(\mathbf{x}))_{i}\in(\operatorname{Sym}_{d}(\mathbb{R}^{n}))^{m}$. For this purpose we will sample $N$ input vectors $\hat{\mathbf{x}}_{1},\dots,\hat{\mathbf{x}}_{N}\in\mathbb{R}^{n}$ according to some probability distribution of our choice. We evaluate $(f_{i}(\mathbf{x}))_{i}$ at these points to obtain $\hat{\mathbf{y}}_{j}\coloneqq(f_{i}(\hat{\mathbf{x}}_{j}))_{i}\in\mathbb{R}^{m}$ for $j=1,\dots,N$; the pairs $(\hat{\mathbf{x}}_{j},\hat{\mathbf{y}}_{j})_{j}$ constitute the training data.
@@
Therefore, the loss function is an $E^{i}_{\alpha,\beta}$-weighted Euclidean distance in the space $(\operatorname{Sym}_{d}(\mathbb{R}^{n}))^{m}$, the ambient space of the neurovariety $\mathcal{V}_{\mathbf{d},r}$. For a general choice of training data $(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})_{i}$, the linear transformation $E$ is general, thus, the learning degree equals the generic Euclidean distance degree of $\mathcal{V}_{\mathbf{d},r}$.
@@
Suppose $\mathbf{x}_{1},\ldots,\mathbf{x}_{N}$ are drawn independently from a fixed distribution $\mathcal{D}$ with known moments $\mu_{k}$, $k=1,2,\ldots$ Based on the law of large numbers, for any $\epsilon>0$,
@@
Here is an explicit result for the architecture $\mathbf{d}=(2,2,k),\leavevmode\nobreak\ r=2$.
@@
The generic ED degree of $\mathcal{V}_{(2,2,k),2}$ is $38k^{2}-12k+$ 3, i.e., the learning degree of the PNN with architecture $(\mathbf{d}=(2,2,k),r=2)$ is $38k^{2}-12k+$ 3.
@@
where $-SS\SS$ and $\mathcal{Q}$ are the universal sub- and quotient bundle of the Grassmannian, respectively, and $i,j=0,1,\dots,m(n-k)$.

### v2

@@ 1. Introduction
In particular, researchers study various activation functions to understand their role in introducing non-linearity, affecting gradient propagation, and influencing computational efficiency. Polynomial activation functions have gained interest for their ability to introduce higher-order interactions between inputs, allowing networks to model complex, non-linear phenomena more effectively [OPP03]. Although feedforward neural networks with polynomial activation functions are well-known to be non-universal approximators [HSW89], there are many practical tasks where architectures with polynomial activations have outperformed other ones, especially in environments where data relationships are polynomial in nature. In particular, polynomial neural networks have led to state-of-the-art results in engineering tasks like face detection from cluttered images [HSHK03], image generation [CMB+20], 3D shape recognition [YHN+21], as well as financial applications like forecasting trading signals [GHL11], uncertain natural frequency quantification [DNM+16], and estimating stock closing indices [NM18]. However, the choice of degree and the potential for overfitting are critical considerations in applying polynomial neural networks to individual practical tasks.
@@
In this paper, we perform an algebro-geometric study of neuromanifolds [^1] By an abuse of terminology we use the term neuromanifold to denote the functional space of a neural network, although this space is almost never a smooth manifold. and their Zariski closures, neurovarieties, following the previous work by [KTB19]. In §2, we give the background on polynomial neural networks, neuromanifolds and neurovarieties. Neuromanifolds provide a natural generalization of the set of symmetric tensors of bounded symmetric rank. In §3, we recall the connection between homogeneous polynomials and symmetric tensor decompositions, and characterize neuromanifolds for some shallow polynomial neural networks. In §4, we describe different approaches for studying neurovarieties. The dimension of the neurovariety provides a measure for the expressivity of the network. In §5, we study these dimensions. For a shallow polynomial neural network with a single output unit, this corresponds to the Alexander–Hirschowitz Theorem in neural network settings. In the deep case, we present conjectures supported by empirical evidence. Finally, in §6, we study the optimization process of a polynomial neural network: from a static perspective, we describe the complexity of the optimization landscape by its learning degree and compute it for a family of architectures; for the dynamic optimization process, we review the backpropagation algorithm and summarize known results on linear neural networks from previous literature in both the machine learning and algebraic statistics communities.
@@
Main contributions. For those new to the topic, we recommend coming back to this paragraph on the main contributions after reading Section 2. We characterize the neuromanifold $\mathcal{M}_{(d_{0},d_{1},1),2}$ in Lemma 3.3, $\mathcal{M}_{(d_{0},1,d_{2}),2}$ in Lemma 3.4, $\mathcal{M}_{(d_{0},1,d_{2}),r}$ in Lemma 3.6, and $\mathcal{M}_{(d_{0},d_{1},d_{2}),r}$ with $d_{1}\geq\binom{r+d_{0}-1}{r}$ in Proposition 3.7. We characterize the neurovariety $\mathcal{V}_{(2,2,d_{2}),2}$ in Proposition 4.1 and partially the neuromanifold $\mathcal{M}_{(2,2,d_{2}),2}$ in Proposition 4.2 by studying the fibers of the PNN, the neurovariety $\mathcal{V}_{(3,3,3),2}$ in Example 4.4 by using Grassmannians and the neurovariety $\mathcal{V}_{(2,2,2,2),2}$ in Example 4.6 by applying the Hilbert–Burch Theorem. In Proposition 5.4, we show that if a PNN has a layer of width one, then the widths of layers after the width-$1$ layer except for the output layer do not influence the neuromanifold. In Corollary 5.9, we show that if a PNN has a defective subnetwork, then it is itself defective unless it is filling. In Section 5.2, we describe a family of non-trivial symmetries for the neuromanifold $\mathcal{V}_{(5,7,1),3}$ that does not have expected dimension. We introduce the learning degree, a characteristic of the optimization landscape, in Definition 6.1. In Theorem 6.3, we show that learning degree of a PNN with respect to the Euclidean loss function exists and it is bounded above by the generic Euclidean distance degree of its neurovariety. In Theorem 6.5, we prove that the learning degree of $\mathcal{V}_{(2,2,k),2}$ for $k\geq 2$ is at most $38k^{2}-12k+$ 3.
Acknowledgments. This paper was written in connection with the “Apprenticeship Week: Varieties from Statistics” which took place at IMSI as part of the long program “Algebraic Statistics and Our Changing World”. We would like to thank Miles Bakenhus and Maksym Zubkov who participated in the work of this group part of the time. Lemma 3.5 is due to Maksym Zubkov. We would like to thank Elizabeth Gross, Leonie Kayser, Joe Kileel, Mark Kong, Guido Montúfar, Alessandro Oneto, Bernd Sturmfels, and Matthew Trager for useful discussions and various suggestions. We thank Lisa Seccia and Nils Sturma for careful reviews of an earlier version of this paper. We also thank Giovanni Luca Marchetti and Vahid Shahverdi for pointing out a mistake in Section 6.1 in a previous version of this article. Part of this research was performed while the authors were visiting the Institute for Mathematical and Statistical Innovation (IMSI), which is supported by the National Science Foundation (Grant No. DMS-1929348).
@@ 2. Background
Figure 1. A neural network architecture with widths $\mathbf{d}=(2,3,3,1)$, input $\mathbf{x}=(x_{1},x_{2})^{T}$ and output $y_{1}$.
Pictorially, a neural network architecture can be represented as in Figure 1. Each node is called a neuron and each column of neurons forms a layer of the network. The first layer is called input layer, the last layer is called output layer and the remaining layers are referred to as hidden layers. A neural network with exactly one hidden layer is called shallow. The number of neurons in the $i^{\text{th}}$ layer is the $i^{\text{th}}$ width, denoted $d_{i}$. The vector of widths $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ together with a choice of activation functions $\bm{\sigma}=(\sigma_{1},\sigma_{2},\dots,\sigma_{L-1})$ constitute the architecture of the network.Arrows between layers represent the maps $f_{l}$. Note that all architectures in this paper are considered to be fully-connected feedforward neural networks.
@@
The image of $\Psi_{\mathbf{d},r}$ is the neuromanifold $\mathcal{M}_{\mathbf{d},r}$. Its Zariski closure is the neurovariety $\mathcal{V}_{\mathbf{d},r}$, an affine variety inside $(\operatorname{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$.
The neuromanifold $\Psi_{\mathbf{d},r}$ is a semialgebraic set inside $(\operatorname{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$, which means that it is a finite union of sets that can be defined by polynomial equalities and inequalities. The observation that the neuromanifold $\Psi_{\mathbf{d},r}$ is a semialgebraic set follows from the Tarski–Seidenberg theorem which states that a polynomial image of a semialgebraic set is again semialgebraic.
@@
Let $\mathbf{d}=(d_{0},\dots,d_{i},\dots,d_{L})$ and let $\mathbf{d}^{\prime}=(d_{0},\dots,d^{\prime}_{i},\dots,d_{L})$ be a tuple which differs from $\mathbf{d}$ precisely in the $i^{\text{th}}$ entry for $0<i<L$ and assume $d^{\prime}_{i}\geq d_{i}$. Then $\mathcal{M}_{\mathbf{d},r}\subseteq\mathcal{M}_{\mathbf{d}^{\prime},r}$, in particular $\mathcal{V}_{\mathbf{d},r}\subseteq\mathcal{V}_{\mathbf{d}^{\prime},r}$.
@@
Most ideals of neurovarieties in Table 1 are described by results in the upcoming sections. The ideals for all but two cases ($\mathbf{d}=(3,2,3)$ and $\mathbf{d}=(3,3,3)$) in the table are determinantal. In general, we expect most neurovarieties to be non-determinantal. Among the table, only the neurovariety for the architecture $\mathbf{d}=(3,2,1)$ does not have expected dimension. The following example describes the cases that are not explained by any results in the rest of the paper.
@@ 3.3. Neuromanifolds for $r\in\mathbb{N}$
Let $\mathbb{k}$ be $\mathbb{R}$ or $\mathbb{C}$ and let
[…]
 of
T is k
[…]
 of
[…]
If $S^{n}_{d,k}(\mathbb{k})\subseteq\operatorname{Sym}_{d}(\mathbb{k}^{n})$ has non-empty interior with respect to the Euclidean topology, $k$ is called a typical symmetric rank for order-$d$ tensors of dimension $n$.
Over the complex numbers, there is a unique typical symmetric rank which is called the generic symmetric rank. However, over the real numbers there can be multiple typical symmetric ranks. In the following we will write $S^{n}_{d,k}$ for $S^{n}_{d,k}(\mathbb{R})$. A neuromanifold $\mathcal{M}_{\mathbf{d},r}$ having non-empty interior with respect to the Euclidean topology implies that the architecture $(\mathbf{d},r)$ is filling. The importance of typical symmetric ranks for PNNs comes from the following fact which is a consequence of the discussions above and the inclusion $S^{n}_{d,k}\subseteq\mathcal{M}_{(n,k,1),d}$.
@@
The algebraic boundary $\partial_{\text{alg}}(\mathcal{M})$ of a set $\mathcal{M}$ is the Zariski closure of its topological boundary $\partial(\mathcal{M})$. Little is known about the algebraic boundaries in the cases considered above. Theorem 4.1 in [MMSV16] implies that the algebraic boundary $\partial_{\text{alg}}(\mathcal{M}_{(3,6,1),4})$ has an irreducible component of degree 51.
@@ 4. Neurovarieties
where $M_{i,j}(C)$ is the $2\times 2$ minor of $C$ obtained by taking the determinant of the $i^{\text{th}}$ and $j^{\text{th}}$ column. In particular, the algebraic boundary of $\mathcal{M}_{(2,2,2),2}$ is given by
$\partial_{\text{alg}}\mathcal{M}_{(2,2,2),2}=\mathcal{V}(M_{1,3}(C)^{2}-M_{1,2}(C)M_{2,3}(C)).$
@@
[…]
@@ 5.1. A plethora of conjectures
It is conjectured that the neurovariety attains the expected dimension for large activation degree. Compare this for example with the Alexander–Hirschowitz Theorem which tells us there exist no defective single-output two-layer neurovarieties with activation degree $r\geq 5$. The conjecture appeared as a theorem in [KTB19], but a mistake in the proof was pointed out by Theo Elenius in his BSc thesis. After the present article first appeared as a preprint on arXiv, the conjecture has been proven in [FRWY24, Theorem 12].
@@ 6. Optimization
To minimize empirical risk, gradient-based optimization methods such as gradient descent (GD), stochastic gradient descent (SGD), and adaptive moment estimation (Adam) are commonly used. For convex objective functions, gradient-based algorithms can converge to the unique global minimum with appropriate hyperparameter choices. However, when optimizing neural networks, the loss landscape is typically non-convex. As a result, there is generally no theoretical guarantee that gradient-based methods will reach a global optimum $f^{*}$ for feedforward neural networks.
Recent research on convergence properties in gradient-based algorithms often focuses on the dynamics of the optimization process. One approach analyzes convergence rates by relaxing assumptions on the objective function. For example, it has been shown that for objective functions with a unique global minimum, gradient descent converges to the global minimum at a geometric rate if the function satisfies conditions such as the Polyak–Lojasiewicz inequality [KNS16, LO19, KR20, MGD+21], weak quasi-convexity [HMR18], or the restricted Secant inequality [HMR18, YZY+20, GEIGM22]. Another line of work explores algorithmic regularization. Empirical evidence suggests that large-scale neural networks used in practice are highly over-parameterized, often containing far more trainable parameters than training examples. Consequently, optimization objectives for such high-capacity models tend to have many global minima that interpolate the data perfectly. The choice of optimization algorithm thus introduces an implicit inductive bias, guiding the solution toward specific types of global minima. One of the first results in this direction shows that for almost all linearly separable datasets, gradient descent with any initialization and any bounded stepsize converge in direction to maximum margin separator with unit ${l}_{2}$ norm [SHN+18]. Both approaches build on the assumption of the existence and separation of global minima, while the global picture of the distribution of critical points is not well-understood. Through algebro-geometric approach, we study the geometry of the optimization landscape and provide the first explicit upper bound to the number of critical points when optimizing over a special family of polynomial neural networks.
@@ 6.1. The Learning Degree of a PNN
Let us consider the case of PNNs again. Often one tries to solve the optimization problem (6.1) using gradient based methods, see §6.2. Therefore, one is interested in computing the critical points on the neurovariety of the loss function $\ell_{p_{\mathbf{w}}}$ as defined in (6.2) for a PNN $p_{\mathbf{w}}$. In particular, the number of such points provides an upper bound for the number of functions the optimization process can converge to when using gradient descent.
Definition 6.1.
If there is a finite number of critical points of $\ell_{p_{\mathbf{w}}}$ on the regular locus of $\mathcal{V}_{\mathbf{d},r}$ which is constant for a general choice of training data $(\mathbf{x}_{i},\mathbf{y}_{i})_{i}$, we call this number the learning degree of the network $p_{\mathbf{w}}$ with respect to loss $\ell$, and denote it by $\operatorname{Ldeg}_{\ell}(p_{\mathbf{w}})$.
In the following we are studying the case of the Euclidean loss function. We show that in this case the learning degree is independent of the choice of training data $(\mathbf{x}_{i},\mathbf{y}_{i})_{i}$. An upper bound of the learning degree is provided by the generic Euclidean distance (ED) degree. We would like to thank Matthew Trager for pointing out the connection to the generic ED degree.
@@
Let $(p_{\mathbf{w}}^{(i)}(\mathbf{x}))_{i}\in(\operatorname{Sym}_{d}(\mathbb{R}^{n}))^{m}$ be a tuple of polynomials in $n$ variables $\mathbf{x}=(x_{1},\dots,x_{n})$ occurring as the output of a fixed PNN with weight vector $\mathbf{w}$. Suppose we want the PNN to learn the tuple of polynomials $(f_{i}(\mathbf{x}))_{i}\in(\operatorname{Sym}_{d}(\mathbb{R}^{n}))^{m}$. To this end, we sample $N$ input vectors $\hat{\mathbf{x}}_{1},\dots,\hat{\mathbf{x}}_{N}\in\mathbb{R}^{n}$ according to some probability distribution. We evaluate $(f_{i}(\mathbf{x}))_{i}$ at these points to obtain $\hat{\mathbf{y}}_{j}\coloneqq(f_{i}(\hat{\mathbf{x}}_{j}))_{i}\in\mathbb{R}^{m}$ for $j=1,\dots,N$; the pairs $(\hat{\mathbf{x}}_{j},\hat{\mathbf{y}}_{j})_{j}$ constitute the training data.
@@
Therefore, the loss function is an $E^{i}_{\alpha,\beta}$-weighted Euclidean distance in the space $(\operatorname{Sym}_{d}(\mathbb{R}^{n}))^{m}$, the ambient space of the neurovariety $\mathcal{V}_{\mathbf{d},r}$. For a general choice of training data $(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})_{i}$, each block $E^{i}$ of the linear transformation is generic. Note, however, that all blocks $E^{i}$, for $i=1,\dots,m$, will be equal, hence, the transformation $E$ might be non-generic. We summarize the discussion in the theorem below.
Theorem 6.3.
The learning degree $\operatorname{Ldeg}_{\ell_{2}}(p_{\mathbf{w}})$ of a PNN $p_{\mathbf{w}}$ with architecture $(\mathbf{d},r)$ with respect to the Euclidean loss function exists. It is at most the generic Euclidean distance degree of its neurovariety, i.e.,
$\operatorname{Ldeg}_{\ell_{2}}(p_{\mathbf{w}})\leq\mathrm{EDdeg}_{\mathrm{gen}}(\mathcal{V}_{\mathbf{d},r}).$ (6.4)
If $p_{\mathbf{w}}$ has only a single output neuron, one has equality in (6.4).
@@
Suppose $\mathbf{x}_{1},\ldots,\mathbf{x}_{N}$ are drawn independently from a fixed distribution $\mathcal{D}$ with known moments $\mu_{k}$, $k=1,2,\ldots$ By the law of large numbers, for any $\epsilon>0$,
@@
For the architecture $\mathbf{d}=(2,2,k),\leavevmode\nobreak\ r=2$, we compute the bound in (6.4) explicitly.
@@
For $k\geq 2$, the generic ED degree of $\mathcal{V}_{(2,2,k),2}$ is $38k^{2}-12k+$ 3. Hence, the learning degree of the PNN with architecture $(\mathbf{d}=(2,2,k),r=2)$ is at most $38k^{2}-12k+$ 3.
@@
where $SS\SS$ and $\mathcal{Q}$ are the universal sub- and quotient bundle of the Grassmannian, respectively, and $i,j=0,1,\dots,m(n-k)$.
@@
Remark 6.9.
For the architecture $(\mathbf{d}=(2,2,3),r=2)$, Theorem 6.8 yields $\operatorname{Ldeg}_{\ell_{2}}(p_{\mathbf{w}})\leq 39$. A numerical computation for a random choice of block $E^{i}$ in the linear transformation $E$ suggests that the actual learning degree is $\operatorname{Ldeg}_{\ell_{2}}(p_{\mathbf{w}})\leq 3$. Moreover, we observe that all critical points of the loss function are actually real. This phenomenon should be further studied in future work.